If a to the power of x =b to the power of y= c to the power of z and b square=AC ,then y is equals to

Grade:8

1 Answers

Arun
6 years ago
Dear student
Given ax = by = cz and b2 = ac
Let ax = by = c= k
Consider,  ax = k  ⇒ a = k1/x
Similarly, we get b = k1/y and c = k1/z
Now consider, b2 = ac
⇒ (k1/y)2 = (k1/x)(k1/z)
⇒ k2/y =  k1/x + 1/z
Comparing both the sides, we get
(2/y) = (1/x) + (1/z)
⇒ (2/y) = (x + z)/xz
∴ y = (2xz)/(x + z)
Regards
Arun